OpenStudy (anonymous):

Find dy\dx if y=(x+1)^x

OpenStudy (anonymous):

\[y=(x+1)^x\]\[lny = xln(x+1)\]Diff. both sides w.r.t. x

OpenStudy (anonymous):

plz do it for me

OpenStudy (anonymous):

i have tried but failed in getting answer

OpenStudy (anonymous):

Can you show your work?

OpenStudy (anonymous):

i have a few questions similar to it so explain also

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

did u use this?

OpenStudy (anonymous):

\[y=(x+1)^x\]\[\ln y=x\ln(x+1)\]Diff. both sides w.r.t. x \[\frac{1}{y} \frac{dy}{dx} = x\frac{d}{dx}(\ln (x+1))+ ln(x+1)\frac{d}{dx}x\] Yes, product rule.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

thanks all

OpenStudy (anonymous):

@RolyPoly

OpenStudy (anonymous):

\[\frac{1}{y} \frac{dy}{dx} = x\frac{d}{dx}(\ln (x+1))+ ln(x+1)\frac{d}{dx}x\]\[\frac{1}{y} \frac{dy}{dx} = x(\frac{1}{x+1})+ ln(x+1)(1)\]\[\frac{1}{y} \frac{dy}{dx} = (\frac{x}{x+1})+ ln(x+1)\]\[\frac{dy}{dx} = y[(\frac{x}{x+1})+ ln(x+1)]\]Then sub y= (x+1)^x into dy/dx